A borrower agrees to repay a loan of $3000 by 15 annual repayments of $500, the first repayment being due after five years. Find the annual yield for this transaction.
1
Expert's answer
2020-09-29T14:50:26-0400

Assume that the annual rate is "i". Then we have:

"500(\frac{1}{(1+i)^5}+\frac{1}{(1+i)^6}+..+\frac{1}{(1+i)^{20}})=3000" .

After simplifications we get:

"\frac{500}{(1+i)^5}\frac{(1-\frac{1}{(1+i)^{16}})}{1-\frac{1}{(1+i)}}=3000"

From it we receive:

"{(\frac{1}{(1+i)^5}-\frac{1}{(1+i)^{21}})}=6(1-\frac{1}{(1+i)})"

Solving the latter numerically, we obtain:

"\frac{1}{(1+i)}=0.91903"

From the latter we obtain i:

"i\approx0.0881."

Answer: 0.0881(the values is rounded to 4 decimal places)